# InTAR: Inter-Task Auto-Reconfigurable Accelerator Design for High Data Volume Variation in DNNs

### v1

@@ IV-B Designing Reconfigurable Modules: A Case Study in HLS
Following the architecture template, there are three types of reconfigurations: data movement, compute, and control. All of them can be implemented in HLS efficiently with conditional dataflow, which utilizes if-else statements to declare the behavior changes and exploit HLS’s resource binding pass to merge interconnects and logic as much as possible.
@@
Compute Reconfiguration: Based on the definition of the architecture template, the reconfigurable PEA can change the input operands and the precision. Figure 4 illustrates a PEA switching the data source between stages, where each source is interpreted as a different precision. To change the input operands, we adopt a similar method of data movement reconfiguration to extract inputs from two source buffers conditioned on the stage index. Inputs are packed for data parallelism. Then, when computations start, each PE will select part of the input operand and pad with zeros to make sure the bitwidths are uniform across stages. This guides the HLS to bind compute resources between stages. The two steps cannot be merged, otherwise the HLS will instantiate two PEAs with different precision specifications.
@@
Figure 5: Example of control reconfigurations in HLS. Left: loop bound control. Right: data-dependent conditional dataflow.
All accelerators in the InTAR paradigm are covered by the compositions of these three types of reconfigurations, indicating complete support in HLS.
@@ IV-C Important Considerations in Placement and Routing
In order to achieve rapid development, we need to make sure that the generated design is both valid and performant. Following are several important considerations of implementing a place-able and routable InTAR design with high frequency.
@@
On-chip Memory Parallel Access Dimension. To serve multiple MAC units in each PE, scratchpad memory is partitioned based on access patterns. In InTAR, the PEA in the CCs may require different memory access patterns. For example, Figure 6 shows two dependent matrix multiply operations, with the dimension indicating the number of data accessed in parallel. The first tiled matrix multiply will compute 2048 MACs simultaneously ($4\times 8\times 64$) and the second matrix multiply computes 1024 MACs ($4\times 4\times 64$) if we align the parallel access pattern. Since we reuse the PEA across tasks, we need to align the throughput between tasks. Thus, it is necessary to further partition the inputs of the second matrix multiplication by two, causing high-fanout nets (HFN) at the write ports, increasing cycle time, and reducing the design frequency. To address this issue, we suggest tiling the operands such that the output tile has equal lengths at each dimension to align parallel read/write dimensions even after transposing, e.g., having $16\times 16$ output tiles in the example.
@@
TABLE III: Descriptions of multi-task kernels used in evaluation 1 with the corresponding parameters and data size configurations.
Kernel Description DNN Applications Parameters Testbench Data Size Config Data Variation Pattern
Self Attention (Attn) A set of dependent matrix multiplications that extract the contextual relationships between tokens in a sequence. Widely used in sequence modeling (Transformer [35], ViT [21], GAT [37]) 256 token sequences. The hidden dimension is 1024. Min data size: 0.5 MB, Max data size: 1.0 MB, On-chip memory constraint: 0.625 MB Increase $\rightarrow$ Decrease $\rightarrow$ Increase $\rightarrow$ Decrease
FFN Layer (FFN) Sequence of linear projections for feature extractions. Input is a single vector. Elementary block in many CV and LLM applications (ResNet [22], YoloV3 [16], Transformer) Input size = final output size = 256. First layer output size = 1024, second layer output size = 2048. Min data size: 0.5 KB, Max data size: 2 KB, On-chip memory constraint: 1 KB Increase $\rightarrow$ Increase $\rightarrow$ Decrease
Multi-layer CNN (M-CNN) Sequence of convolutions with upsampling and pooling layers Used in many CV applications (ResNet, VGG [36]) Input size = 224, kernel size = 3, $2\times 2$ upsampling and pooling size. Min data size: 98 KB, Max data size: 1.57 MB, On-chip memory constraint: 125 KB. Increase $\rightarrow$ Increase $\rightarrow$ Decrease $\rightarrow$ Decrease
Variational Autoencoder (VAE) Contains an encoder in convolutions for the latent space distribution and a decoder in transpose convolutions. Used for image compression and the generator model in GAN (VAE-GAN [17]) Input size = 28, kernel sizes are 4 and 8, 2 channels and 2 filters Min data size: 1.3 KB, Max data size: 3.1 KB, On-chip memory constraint: 1.5 KB Decrease $\rightarrow$ Decrease $\rightarrow$ Increase $\rightarrow$ Increase
Gating Network (GN) Two parallel linear projections with element-wise product and a sequential linear projection. Llama-family LLMs [34] Sequence length = 32, input dimension = 4096, hidden dimension = 11008 Min data size: 0.25 MB, Max data size: 0.67 MB, On-chip memory constraint: 0.5 MB Increase $\rightarrow$ Decrease
@@ V-A Experiment Setup
To demonstrate the broadness of InTAR in HDV DNN accelerations, we construct a testbench with five multi-task kernels that can serve as a standalone DNN application (Multi-layer CNN, Variational Autoencoder) or participate in part of the DNN applications (Self Attention, FFN Layer, Gating Network). Table III illustrates these kernels in the evaluation. Unlike existing benchmarks [44, 28] for FPGA-based accelerator evaluation, our testbench has three features crucial for HDV: ① it covers various DNN applications in computer vision, language modeling, graph processing, etc., displayed in the ”DNN Applications” column; ② it adopts parameters used in the real applications (e.g., Gating Network has the identical dimensionalities as Llama 2[34]); and ③ kernels are all HDV, and we impose a memory constraint between the minimum and maximum input/output data sizes so that InTAR can exploit its ability to switch execution modes.
Using our testbench, we compare InTAR with human-optimized dataflow and sequential accelerators under the same resource constraints, including on-chip memory size, the number of DSPs, and off-chip memory bandwidth utilization. All designs are implemented utilizing Xilinx Vitis HLS 2021.2 with TAPA [19] and evaluated on Xilinx Alveo U280 FPGA board, with clock frequency at 300 MHz. We chose the human-optimized designs as the baseline since: ① There is a lack of existing designs of the kernels in our testbench with the same model and hardware configurations, and ② existing frameworks [15, 6, 9] for DNN application design generation on FPGAs do not handle off-chip memory communications or do not consider the extra resource constraints introduced to evaluate HDV DNNs.
Figure 7: Speedup and normalized DSP efficiency of InTAR over dataflow and sequential execution for the five multi-task kernels in the testbench.
@@ V-B Analysis
Figure 7 shows the speedup and DSP efficiency of InTAR over sequential and dataflow accelerators. Overall, InTAR achieves a speedup of 7.1×7.1$\times$7.1 × and 1.8×1.8$\times$1.8 × over sequential and dataflow accelerators, along with 7.5×7.5$\times$7.5 × and 1.9×1.9$\times$1.9 × higher DSP efficiency (GOP/s/DSP) in geometric mean. The performance boost of InTAR over sequential accelerators mainly comes from the reduction of off-chip memory access, where InTAR has 76% lower off-chip memory access volume than the sequential accelerators in geomean over the five kernels. For dataflow accelerators, InTAR is significantly faster in self-attention and multi-layer CNN than the other three kernels. There are two explanations. First, the two kernels mentioned have a higher data reuse between dependent tasks for matrix multiplications and convolutions, which requires data buffering and preprocessing before continuing the computations (e.g., the pooling layer of the multi-layer CNN will wait until producing two rows of outputs from the previous layer to downsample the data). Since the selected kernels have a long chain of dependent tasks, pipeline stall latency may dominate the end-to-end latency. Second, InTAR can better reuse off-chip memory ports for these two kernels. For example, after loading inputs and finishing the first convolution layer in the multi-layer CNN, InTAR can utilize the memory ports allocated to the inputs to write outputs. Thus, InTAR has a higher effective off-chip read/write throughput than dataflow accelerators which only use the dedicated memory port for input/output. For Gating Networks, InTAR cannot further reduce latency compared with dataflow accelerators since only input and output data sizes are within the memory constraint and our schedule for InTAR is identical to pipelining all tasks.
@@ VI Evaluation 2: GPT-2 Input Prefilling
Figure 8: A sample design of InTAR mapped to Versal VPK180 FPGA for GPT-2 Medium model. (a) Overall architecture and parameters. Each SLR contains two CCs and one SFU. (b) The scheduled timeline. The color of the blocks corresponds to the PEs or SFUs shown in (a).
@@ VI-A Experiment Setup
We choose the GPT-2 medium input prefilling stage as the example of a complex DNN application to further evaluate InTAR and compare it with the corresponding SoTA accelerators. Figure 8(a) illustrates the architecture design for the GPT-2 medium model on the AMD Xilinx Versal VPK180 FPGA, based on the template in Section 2. The design has four rows of CCs spreading out to four SLRs of the FPGA. Each SLR includes two CCs and an SFU. Figure 8(b) depicts the GPT-2 medium schedule. Thanks to the grid structure of InTAR’s architecture template, we can adapt the design conveniently to other FPGA boards with slight modifications. We also implemented an InTAR design example for Alveo U280 FPGA, which contains three SLRs.
To generate the circuit design, we utilize Xilinx HLS with the TAPA framework [19] and prototype with the Vitis 2021.2 toolchain for U280 and Vitis 2024.1 for VPK180. The floorplanning of the CCs and SFUs is predetermined based on the architecture template and the rest of the modules (e.g., auxiliary buffers) are done by Autobridge [20]. The designs are both placed and routed for U280 and VPK180. We select Allo [9] and DFX [23] as the SoTA accelerators for the GPT-2 medium model to compare their throughput with InTAR.
@@
Max Power 75 W 180 W 70 W 250 W 300 W
Max Perf. 8.09 TOP/s 20.7 TOP/s 65.13 TOP/s 311.84 TOP/s 181 TOP/s
@@ VI-B Analysis
Comparison with FPGA-based Accelerators. Figure 9 presents the latency, DSP efficiency, and power efficiency comparison between FPGA and GPU accelerators. Compared to Allo and DFX, InTAR on U280 is 7.99×7.$9$ and 14.64×14.64$\times$14.64 × faster, and 2.19×2.19$\times$2.19 × and 3.98×3.98$\times$3.98 × more DSP efficient, respectively. InTAR on VPK180 is 21.39×21.39$\times$21.39 × and 39.14×39.14$\times$39.14 × faster, and 5.66×5.$6$ and 10.44×10.$4$ more DSP efficient than Allo and DFX. InTAR on VPK180 is more DSP efficient than U280 since DSP58 has a higher throughput than DSP48. Additionally, implementing DSPs in dot-product modules on VPK180 trims the logic and further reduces cycles, leading to more than 2×2$\times$2 × boost in DSP efficiency.
Comparison with GPUs. As shown in Figure 9, InTAR on U280 achieves a 1.07×1.07$\times$1.07 × speedup in geometric mean over T4 and 2.41×2.41$\times$2.41 × speedup on VPK180. Especially for short sequences ($L\leq 256$), InTAR’s performance surpasses both T4 and A100 GPUs.
@@ VII Conclusion and Future Work
While we are motivated and focused on DNNs, InTAR can potentially improve the performance of non-DNN HDV applications, which we plan to explore. Furthermore, future work including developing an optimized DSE engine to replace the heuristic-based scheduling and integrating InTAR into compilers are necessary to address InTAR’s limitations.

### v2

@@ IV-B Designing Reconfigurable Modules: A Case Study in HLS
In this work, we require the designer to manually write the reconfiguration designs based on the static schedule. Following the architecture template, there are three types of reconfigurations: data movement, compute, and control. All of them can be implemented in HLS efficiently with conditional dataflow, which utilizes if-else statements to declare the behavior changes and exploit HLS’s resource binding pass to merge interconnects and logic as much as possible.
@@
Compute Reconfiguration: Based on the definition of the architecture template, the reconfigurable PEA can change the input operands and the precision. Figure 4 illustrates a PEA switching the data source between stages, where each source is interpreted as a different precision. To change the input operands, we adopt a similar method of data movement reconfiguration to extract inputs from two source buffers conditioned on the stage index. Inputs are packed for data parallelism. Then, when computations start, each PE will select part of the input operand and pad with zeros to make sure the bitwidths are uniform across stages. This guides the HLS to bind compute resources between stages. The two steps cannot be merged, otherwise, the HLS will instantiate two PEAs with different precision specifications.
@@
Figure 5: Example of control reconfigurations in HLS. Left: loop-bound control. Right: data-dependent conditional dataflow.
All accelerators in the InTAR paradigm are covered by the compositions of the three types of reconfigurations mentioned above, indicating complete support in HLS.
@@ IV-C Important Considerations in Placement and Routing
In order to achieve rapid development, we need to make sure that the generated design is both valid and performant. Following are several important considerations of implementing a placeable and routable InTAR design with high frequency.
@@
Matching On-chip Memory Parallel Access Dimension. To serve multiple MAC units in each PEA, scratchpad memory is partitioned by access patterns. In InTAR, the PEA in the CCs may require varying patterns between tasks. For example, Figure 6 shows two dependent GEMMs: the first computes 2048 MACs simultaneously ($4\times 8\times 64$), and the second computes 1024 MACs ($4\times 4\times 64$). Since the PEA is reused across tasks, we aim to equalize the throughputs of the two matrix multiplies for efficiency, necessitating further partitioning of the second GEMM. This causes high-fanout nets at the write ports, increases cycle time, and reduces frequency. We address this by tiling operands so that output tiles have equal dimensions (e.g., $16\times 16$), aligning parallel read/write dimensions even after matrix transpose.
@@
TABLE III: Descriptions of multi-task kernels with the corresponding parameters and data size constraints. The on-chip memory constraints are artificial to make the kernel HDV.
Kernel Description DNN Applications Parameters Data Size Constraints Data Variation Pattern
Self Attention (Attn) A set of dependent matrix multiplications that extract the contextual relationships between tokens in a sequence. Widely used in sequence modeling (Transformer [5], ViT [36], GAT [37]) 256 token sequences. The hidden dimension is 1024. Min data size: 0.5 MB, Max data size: 1.0 MB, On-chip memory constraint: 0.625 MB Increase ($Q,K$) $\rightarrow$ Decrease ($Q,K\rightarrow A$) $\rightarrow$ Increase ($V$$\rightarrow$ Decrease ($A,V\rightarrow V^{\prime}$)
FFN Layer (FFN) Three layers of linear projections for feature extractions. Input is a single vector. Elementary block in many CV and LLM applications (ResNet [15], YoloV3 [38], Transformer) Input size = final output size = 256. First layer output size = 1024, second layer output size = 2048. Min data size: 0.5 KB, Max data size: 2 KB, On-chip memory constraint: 1 KB Increase $\rightarrow$ Increase $\rightarrow$ Decrease
Multi-layer CNN (M-CNN) Four layers of convolutions with upsampling and pooling layers Used in many CV applications (ResNet, VGG [39]) Input size = 224, kernel size = 3, $2\times 2$ upsampling and pooling size. Min data size: 98 KB, Max data size: 1.57 MB, On-chip memory constraint: 125 KB. Increase $\rightarrow$ Increase $\rightarrow$ Decrease $\rightarrow$ Decrease
Variational Autoencoder (VAE) Contains an encoder in convolutions for the latent space distribution and a decoder in transpose convolutions. Used for image compression and the generator model in GAN (VAE-GAN [40]) Input size = 28, kernel sizes are 4 and 8, 2 channels and 2 filters Min data size: 1.3 KB, Max data size: 3.1 KB, On-chip memory constraint: 1.5 KB Decrease (encoder 1) $\rightarrow$ Decrease (encoder 2) $\rightarrow$ Increase (decoder 1$\rightarrow$ Increase (decoder 2)
Gating Network (GN) Two parallel linear projections with element-wise product and a sequential linear projection. Llama-family LLMs [3] Sequence length = 32, input dimension = 4096, hidden dimension = 11008 Min data size: 0.25 MB, Max data size: 0.67 MB, On-chip memory constraint: 0.5 MB Increase (up projections) $\rightarrow$ Decrease (element-wise product, down projection)
@@ V-A Experiment Setup
To demonstrate the broadness of InTAR in HDV DNN accelerations, we construct a testbench with five multi-task kernels that can serve as a standalone DNN application (Multi-layer CNN, Variational Autoencoder) or participate in part of the DNN applications (Self Attention, FFN Layer, Gating Network). Table III illustrates these kernels in the evaluation. Unlike existing benchmarks [41, 42] for FPGA-based accelerator evaluation, our testbench has three features crucial for HDV: ① it covers various DNN applications in computer vision, language modeling, graph processing, etc., displayed in the ”DNN Applications” column; ② it adopts hyperparameters used in the real applications (e.g., Gating Network has the identical dimensionalities as Llama 2[3]); and ③ kernels are all HDV, and we impose a memory constraint between the minimum and maximum input/output data sizes so that InTAR can exploit its ability to switch execution modes.
Using our testbench, we compare InTAR with human-optimized dataflow and sequential accelerators, each designed in four weeks. All designs are evaluated under identical resource constraints, including on-chip memory capacity, the number of DSPs, and off-chip memory bandwidth utilization. All designs are implemented utilizing Xilinx Vitis HLS 2021.2 with TAPA [43] and evaluated on Alveo U280 FPGA board, with a clock frequency of 300 MHz. We chose the human-optimized designs as the baseline since: ① There is a lack of existing designs of the kernels in our testbench with the same model and hardware configurations, and ② existing frameworks [44, 7, 6] for DNN application design generation on FPGAs do not handle off-chip memory communications or do not consider the extra resource constraints introduced to evaluate HDV DNNs.
Figure 7: Speedup and DSP efficiency of InTAR over dataflow and sequential execution for the five multi-task kernels in the testbench. Values are normalized from sequential accelerators.
@@ V-B Analysis
Figure 7 shows the speedup and DSP efficiency of InTAR over sequential and dataflow accelerators. Overall, InTAR achieves a speedup of 7.1×7.1$\times$7.1 × and 1.8×1.8$\times$1.8 × over sequential and dataflow accelerators, along with 7.5×7.5$\times$7.5 × and 1.9×1.9$\times$1.9 × higher DSP efficiency (GOP/s/DSP) in geometric mean. The performance boost of InTAR over sequential accelerators mainly comes from the reduction of off-chip memory access, where InTAR has 76% lower off-chip memory access volume than the sequential accelerators in geomean over the five kernels. For dataflow accelerators, InTAR is significantly faster in self-attention and multi-layer CNN than the other three kernels. There are two explanations. First, the two kernels mentioned have a higher data reuse between dependent tasks for matrix multiplications and convolutions, which requires data buffering and preprocessing before continuing the computations (e.g., the pooling layer of the multi-layer CNN will wait until producing two rows of outputs from the previous layer to downsample the data). Since the selected kernels have a long chain of dependent tasks, pipeline stall latency may dominate the end-to-end latency. Second, InTAR can better reuse off-chip memory ports for these two kernels. For example, after loading inputs and finishing the first convolution layer in the multi-layer CNN, InTAR can utilize the memory ports allocated to the inputs to write outputs. Thus, InTAR has a higher effective off-chip read/write throughput than dataflow accelerators, which only use the dedicated memory port for input/output. For Gating Networks, InTAR cannot further reduce latency compared with dataflow accelerators since only input and output data sizes are within the memory constraint, and our schedule for InTAR is identical to pipelining all tasks.
@@ VI-A Experiment Setup
We choose the GPT-2 medium input prefilling stage as a complex scenario to further evaluate InTAR, which is a frequently used benchmark by prior SoTA FPGA-based LLM accelerators [6, 23]. We selected two platforms (Versal VPK180 [45] and Alveo U280 [34]), which result in different schedules and architecture template parameters based on their resource constraints. For instance, the design for VPK180 has four rows of CCs spreading out to four SLRs of the FPGA, while the design on U280 has 3 rows of CCs to fit its 3-SLR layout. Each SLR includes two CCs and an SFU for both platforms. Thanks to the grid structure of InTAR’s architecture template, we can adapt the design conveniently to other FPGA boards with only slight modifications.
To generate the circuit design, we employ Xilinx HLS with the TAPA framework [43] and prototype using the Vitis 2021.2 that performs well on older platforms such as U280, and Vitis 2024.1 optimized for VPK180. The floorplanning of the CCs and SFUs is predetermined based on the architectural template, while the remaining modules (e.g., auxiliary buffers) are managed by AutoBridge [46]. The designs are both placed and routed for U280 and VPK180. For comparison, we select Allo [6] and DFX [23] as the SoTA accelerators, both of which utilize U280 as the compute platform; accordingly, we include InTAR implemented on U280.
@@
Peak Power 75 W 180 W 70 W 250 W 300 W
Peak Perf. 8.09 TOP/s 20.7 TOP/s 65.13 TOP/s 311.84 TOP/s 181 TOP/s
@@ VI-B Analysis
Comparison with FPGA-based Accelerators. Figure 8 presents the latency, DSP efficiency, and power efficiency comparison between FPGA and GPU accelerators. Compared to Allo and DFX, InTAR on U280 is 7.99×7.$9$ and 14.64×14.64$\times$14.64 × faster, and 2.19×2.19$\times$2.19 × and 3.98×3.98$\times$3.98 × more DSP efficient, respectively. InTAR on VPK180 is 21.39×21.39$\times$21.39 × and 39.14×39.14$\times$39.14 × faster, and 5.66×5.$6$ and 10.44×10.$4$ more DSP efficient than Allo and DFX. InTAR on VPK180 is more DSP efficient than U280 since DSP58 has a higher throughput than DSP48. Additionally, implementing DSPs in dot-product modules on VPK180 trims the logic and further reduces cycles, leading to more than 2×2$\times$2 × boost in DSP efficiency. Furthermore, similar to Evaluation 1 in Section V, we compared InTAR with our manually implemented dataflow and sequential accelerators, achieving 1.8×1.8$\times$1.8 × and 8.3×8.3$\times$8.3 × speedup and 1.7×1.7$\times$1.7 × and 8.2×8.2$\times$8.2 × DSP efficiency boost.
@@ VII Conclusion and Future Work
Although focused on DNNs, InTAR may also enhance non-DNN HDV applications, which we plan to explore. Future work includes developing an optimized DSE engine to replace heuristic scheduling and automate InTAR design compilation.
Acknowledgement
This work was supported in part by PRISM, one of the seven centers in the JUMP 2.0 program sponsored by SRC and DARPA. It is also supported by CDSC industrial partners and the AMD  [^2] J. Cong has a financial interest in AMD HACC Program.